OpenStudy (habbababba1):

Arrange these functions from the greatest to the least value based on the average rate of change in the specified interval.

OpenStudy (habbababba1):

f(x) = x2 + 3x interval: [-2, 3]

OpenStudy (habbababba1):

f(x) = 3x - 8 interval: [4, 5]

OpenStudy (habbababba1):

f(x) = x2 - 2x interval: [-3, 4]

OpenStudy (habbababba1):

f(x) = x2 - 5 interval: [-1, 1]

OpenStudy (habbababba1):

@ganeshie8

OpenStudy (habbababba1):

@mathmate

OpenStudy (habbababba1):

@mattpusey13

OpenStudy (mathmate):

@habbababba1 Hints: step 1: Find the average rate of change of each function within the given domain (=(f(x2)-f(x1)/(x2-x1)) step 2: order the functions in decreasing average rate of change.

OpenStudy (habbababba1):

?